Assume that autonomous consumption is GHS1,791 billion and disposal income is GHS13,000 billion. Using consumption function, compute consumption expenditure if an increase of GHS1,000 in disposable income leads to an increase of GHS720 in consumption expenditure.  


1
Expert's answer
2021-09-14T10:10:59-0400

Consumption expenditure refers to the spending on consumption out of disposable income. When disposable income increases, consumption expenditure also increases, and vice versa. Consumption function (C) comprises autonomous consumption(a), marginal propensity to consume (MPC or b), and disposable income(Yd). Hence, it can be written as-

"C = a + bYd"

It is given, autonomous consumption, a = GHS1,791

                                   Disposable income, Yd = GHS13,000

        Consumption expenditure increases by (?C) = GHS720

          Disposable income increases by (?Yd) = GHS1,000

So, the MPC will be----

"MPC=\frac{\Delta C}{\Delta Yd}\\=\frac{720}{1000}\\=0.72"

Therefore, the consumption expenditure is--

"C=1791+(0.72\times 1300)\\=1791+9360\\=11151"

Hence, the consumption expenditure is GHS 11151